Explain how a linear monoatomic chain behaves as a low pass filter
1
Expert's answer
2020-02-11T09:53:52-0500

The propagation of harmonic waves in a linear chain consisting of equally spaced

masses m connected by linear springs of stiffness k has been studied extensively. The

chain behaves as a low pass filter so that waves can propagate without attenuation

below the frequency "\omega_0=2\sqrt{\frac{k}{m}}" . Above this frequency, the amplitude decays

exponentially and those waves are called evanescent waves. The "[0,\omega_0]" frequency range is called a pass-band and in that range the dispersion relation is given by "\omega=\omega_0 \sin{0.5ka}", where k is the wave number, and a is the distance between two consecutive masses. In that range, waves with different phase velocities "c=\frac{\omega}{k}" and group velocities "c_g=\frac{d\omega}{dk}".

For long waves "(k\to 0)" , the chain behaves as a rod governed by the classical wave

equation. Several higher order continuum models are derived from the dispersion

relation.